2. Calculating marginal revenue from a linear demand curve The blue curve on the following graph represents the demand curve facing a firm that can set its own prices. Use the graph input tool to help you answer the following questions. You will not be graded on any changes you make to this graph Note: Once you enter a value in a white field, the graph and any corresponding amounts in each grey field will change accordingly
Answer #1
Q P TR
0 100 0
10 80 800
20 60 1200
25 50 1250
30 40 1200
40 20 800
50 0 0

1250 1125 Total Revenue TOTAL REVENUE (Dollars) 0 5 10 40 45 50 15 20 25 30 35 QUANTITY (Number of units)

MR of 10th unit = (10*80)-(9*82) = 800-738 =62

MR of the 20th unit = (20*60)-(19*62) = 1200-1178 = 22

Marginal Revenue MARGINAL REVENUE (Dollars) O 5 10 15 35 40 45 50 20 25 30 QUANTITY (Units)When TR is increasing, MR is Positive